OpenStudy (anonymous):

sin x = 3/5 , π/2 < x < π A.cos(x/2) B. sin(x/2) C. Tan(x/2)

OpenStudy (anonymous):

what do we have to find out?

OpenStudy (anonymous):

A. Cos (x/2) B. Sin (x/2) C. Tan (x/2)

OpenStudy (anonymous):

so do you know how to find x? in sin x = 3/5

OpenStudy (anonymous):

@jewest19

OpenStudy (anonymous):

someone can take over. not wasting mi time waiting when i have a geology exam to revise for

OpenStudy (anonymous):

if i knew i wouldnt post it. Dont worry about wasting your time, I was trying to help someone else.

OpenStudy (anonymous):



OpenStudy (anonymous):

use "half angle" formula to find \(\cos(\frac{x}{2})\)

OpenStudy (anonymous):

i.e. use \[\cos(\frac{x}{2})=\sqrt{\frac{1+\cos(x)}{2}}\]

OpenStudy (anonymous):

in this case the answer is positive, and also in this case you can see from the triangle that \(\cos(x)=\frac{4}{5}\)

OpenStudy (anonymous):

4/5 is incorrect, thats what i thought it was

OpenStudy (anonymous):

it is not \(\frac{4}{5}\)

OpenStudy (anonymous):

\(\cos(x)=\frac{4}{5}\) you are looking for \(\cos(\frac{x}{2})\)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

replace \(\cos(x)\) in the "half angle" formula i wrote above by \(\frac{4}{5}\) to find the answer

OpenStudy (anonymous):

put \(\cos(x)=\frac{4}{5}\) here : \[\cos(\frac{x}{2})=\sqrt{\frac{1+\cos(x)}{2}}\]

OpenStudy (anonymous):

ok im working it

OpenStudy (anonymous):

\[\frac{3\sqrt{2}}{ \sqrt{5} }\]

OpenStudy (anonymous):

@satellite73 idk! i got it wrong again

OpenStudy (whpalmer4):

\[\cos(\frac{x}{2}) = \sqrt{\frac{1+\cos(x)}{2}} = \sqrt{\frac{1+\frac{4}{5}}{2}} = \sqrt{\frac{\frac{5}{5}+\frac{4}{5}}{2}} =\sqrt{\frac{\frac{9}{5}}{2}}=\sqrt{\frac{9}{10}} = \]\[\frac{\sqrt{9}}{\sqrt{10}} = \frac{3}{\sqrt{10}} = \frac{3\sqrt{10}}{10}\]